Two point charges are located on the x-axis of a coordinate system: π‘ž_1 =-1.5 𝑛𝐢 is at π‘₯ = +4.0 π‘π‘š, and π‘ž_2=βˆ’3.5 𝑛𝐢 is at π‘₯ = +2.0 π‘π‘š. What is the total electric force exerted by π‘ž_1 and π‘ž_2 on a π‘ž_3= 2.0 𝑛𝐢 at π‘₯ = +3 cm?

1
Expert's answer
2022-04-13T09:20:58-0400

"F=kq_3(\frac{q_1}{x_1^2}+\frac{q_2}{x_2^2})=2~\mu N."